# Gaussian fluctuations for stochastic Volterra equations with small noise

### v1

@@ 1 Introduction
where the initial data $x_{0}\in\mathbb{R},\mbox{ }(B_{t})_{t\in[0,T]}$ is a standard Brownian motion and $b,\sigma$ are measurable functions on $0,T]^{2}\times\mathbb{R}.$
It is well known that the class of stochastic Volterra equations was first studied by Berger and Mizel in [1]. Since then this class has been widely used in the modelling of numerous random phenomena (fluid turbulence [4], DNA patterns [14] and mathematical finance [11], etc). For small values $\varepsilon,$ the equation (1.1) forms a dynamical system with small noise. In this context, the asymptotic behavior of $X_{\varepsilon,t}$ as $\varepsilon\to 0$ has been investigated by various authors. However, most of papers are devoted to the large and moderate deviation results, see e.g. [9, 10, 13, 16] and references therein.
In this paper, our aim is to analyze the Gaussian fluctuation of $X_{\varepsilon,t}$ around its limit. It should be noted that, in the last years, the Gaussian fluctuation for dynamical systems with small noise has been studied by several authors, see e.g. [2, 6, 8] and references therein. We observe that, as $\varepsilon\mbox{ tends to }0,$ the solution $(X_{\varepsilon,t})_{t\in[0,T]}$ of the equation (1.1) converges to a deterministic process $(x_{t})_{t\in[0,T]}$ solving the following integral equation
@@
The results of [15] are qualitative and hence, the next important problem is to provide quantitative estimates for the rate of convergence. Here we use the total variation distance to characterize the rate of convergence:
$d_{TV}(\tilde{X}_{\varepsilon,t},Y_{t}):=\sup_{A}|P(\tilde{X}_{\varepsilon,t}\in A)-P(Y_{t}\in A)|=\frac{1}{2}\sup_{g}|Eg(\tilde{X}_{\varepsilon,t})-Eg(Y_{t})|,$
where the supremum is running over all Borel subsets $A$ of the real line and all measurable functions bounded by $1.$ Then, we use the techniques of Malliavin calculus to obtain an explicit estimate for $d_{TV}(\tilde{X}_{\varepsilon,t},Y_{t}).$ Furthermore, we also show that the rate of convergence is optimal order.

### v2

@@ 1 Introduction
where the initial data $x_{0}\in\mathbb{R},\mbox{ }(B_{t})_{t\in[0,T]}$ is a standard Brownian motion and $b,\sigma$ are measurable functions on $0,T]^{2}\times\mathbb{R}.$ It is well known that the class of stochastic Volterra equations was first studied by Berger and Mizel in [2]. Since then this class has been widely used in the modelling of numerous random phenomena such as fluid turbulence [5], DNA patterns [15] and mathematical finance [1, 12], etc.
For small values $\varepsilon,$ the Volterra equation (1.1) forms a dynamical system with small noise. We recall that the study of asymptotic behaviors of dynamical systems with small noise has a long history beginning in 1970 with the results of Ventzell & Freidlin [17]. In this context, one of the fundamental problems is to study the convergence of the considering system to its deterministic limit. The convergence can be described via large deviation principle, central limit theorem and moderate deviation principle, etc. In fact, the system (1.1), the asymptotic behavior of $X_{\varepsilon,t}$ as $\varepsilon\to 0$ has been investigated by various authors. However, most of papers are devoted to the large and moderate deviation results, see e.g. [10, 11, 14, 18] and references therein.
In this paper, our aim is to analyze the Gaussian fluctuation of $X_{\varepsilon,t}$ around its limit. More specifically, we prove a central limit theorem result with explicit estimates for the rate of convergence. It should be noted that, in the last years, the Gaussian fluctuation for dynamical systems with small noise has been studied by several authors, see e.g. [3, 7, 9] and references therein. We observe that, as $\varepsilon\mbox{ tends to }0,$ the solution $(X_{\varepsilon,t})_{t\in[0,T]}$ of the equation (1.1) converges to a deterministic process $(x_{t})_{t\in[0,T]}$ solving the following integral equation
@@
The results of [16] are qualitative and hence, the next important problem is to provide quantitative estimates for the rate of convergence via certain distances. Among others, the Wasserstein, Kolmogorov and total variation distances have been commonly used in the literature. The the Wasserstein distance
$d_{W}(\tilde{X}_{\varepsilon,t},Y_{t}):=\sup_{|g(x)-g(y)|\leq|x-y|}|Eg(\tilde{X}_{\varepsilon,t})-Eg(Y_{t})|$
is easy to bound by using the estimate (3.16) below. On the other hand, the Kolmogorov distance is dominated by total variation distance. Hence, we will focus on bounding the total variation distance defined by
$d_{TV}(\tilde{X}_{\varepsilon,t},Y_{t}):=\frac{1}{2}\sup_{g}|Eg(\tilde{X}_{\varepsilon,t})-Eg(Y_{t})|,$
where the supremum is running over all measurable functions bounded by $1.$ Our main tools are the techniques of Malliavin calculus. Particularly, we use a general result established in our previous paper [6] to obtain the optimal estimate of the form
$d_{TV}(\tilde{X}_{\varepsilon,t},Y_{t})\sim O(\varepsilon)\,\,\,\text{as}\,\,\varepsilon\to 0.$
We also would like to emphasize that our results are able to apply to the class of stochastic Volterra equations with singular kernels.